find the center (h,k) and radius r of the circle with the given equation
(x-6)^2+y^2=81

Answer by jim_thompson5910(35256)

find the center (h,k) and radius r of the circle with the given equation
(x-6)^2+y^2=81

-------------------------------------------------------------------------------

(x-6)^2+y^2=81

(x-6)^2+(y)^2=81

(x-6)^2+(y-0)^2=81

(x-6)^2+(y-0)^2=9^2


The equation (x-6)^2+(y-0)^2=9^2 is in the form (x-h)^2+(y-k)^2=r^2 where

h = 6
k = 0
r = 9

So the center is (6,0) and the radius is 9